# Through the River: Understanding the Benefit of Schedule-Free Methods for Language Model Training

### v1

@@ 4.4 A Reformulation of Schedule-Free Optimizer
Motivated by Figures 5 and 6, we now examine the dynamics of the ${\mathbf{y}}_{t}$ iterates. Following Defazio et al. (2024) and Morwani et al. (2025), define the momentum variable $\mathbf{m}_{t}\!\coloneqq\!\tfrac{\mathbf{x}_{t}-\mathbf{z}_{t+1}}{\gamma}$. The Schedule-Free update in (SF) is then equivalent to
@@
i.e. SF is simply a momentum-based optimizer update on $\mathbf{y}_{t}$. The full derivation is provided in Section D.2.
@@ 4.5 Central Flow Analysis
As $\beta_{1}$ increases, $\sigma^{2}$ decreases; thus, larger values of $\beta_{1}$ suppress oscillations along the hill directions, keeping the $\mathbf{y}_{t}$ iterates more closely aligned with the river—consistent with the empirical observation in Figure 4. A complete derivation, including the central flow of SF-GD, is provided in Section D.3.
@@ 5 A Refined and Robust Schedule-Free Optimizer
While SF-AdamW achieves strong performance, it is highly sensitive to momentum hyperparameters and degrades under large batch sizes (Zhang et al., 2025; Morwani et al., 2025). Building on the insights from Section 4, we revisit these issues and propose a refined variant that addresses them.
@@
When $\beta$ is small, the weights $\{\alpha_{t}\}$ become stretched, overemphasizing early iterates and preventing $\mathbf{x}_{t}$ from closely tracking the river, as shown in Figure 7. This also explains Figure 2, where applying EWA to ${\mathbf{x}}_{t}$ offers no benefit: since ${\mathbf{x}}_{t}$ is already a weighted average of ${\mathbf{y}}_{t}$, further averaging merely flattens the weights and weakens alignment with the river.
Figure 7: Averaging weights in SF methods. Smaller values of $\beta$ flatten the averaging weights $\{\alpha_{t}\}$.
More fundamentally, $\beta$ plays a dual role: it controls both the momentum update of ${\mathbf{y}}_{t}$ and the width of the averaging window for ${\mathbf{x}}_{t}$. The optimal value for each may differ, and this mismatch can hinder performance. In large-batch training, a narrower window (i.e., larger $\beta$) is preferred, but this also slows ${\mathbf{y}}_{t}$ updates, as both $\beta c_{t+1}$ and $(1-\beta)$ in (SFy) become vanishingly small when $t$ is large.
@@
The full derivation is given in Section D.4, and pseudocode is provided in Algorithm 2. As shown in Figures 13 and 14, the weights $\{\alpha_{t}\}$ closely follow the theoretical approximation across different values of $\beta$, $C$, and $T$.
This modification makes the averaging weights $\{\alpha_{t}\}$ depend solely on $C$, allowing $\beta$ to independently control the momentum on $\mathbf{y}_{t}$. In other words, $C$ decouples the momentum and averaging behavior. Notably, both vanilla and refined SF use an averaging window that grows with $T$, unlike fixed-width schemes such as EWA.
@@
We evaluate the performance of our refinement to SF-AdamW using the experimental setup as in Section 3. Keeping all other hyperparameters fixed from the tuned vanilla configuration and varying only $C$ (with $C=1/(1-\beta_{1})$ recovering the original method), we observe:
[…]
  • Improved best-case performance. In the best vanilla setup ($\beta_{1}=0.95$), setting $C=200$ leads to further reductions in validation loss; similar gains are observed with $\beta_{1}=0.9$ (Figure 8-middle).
  • Large-batch setting. With 2M-token batches, vanilla SF-AdamW ($\beta_{1}=0.98$) lags behind AdamW with cosine schedule, whereas setting $C=200$ matches its final performance (Figure 8-right).
These results show that decoupling momentum and averaging via introducing $C$ eliminates the tradeoff inherent in vanilla SF, yielding a more robust and scalable optimizer.
Figure 8: Refined SF-AdamW. Left: Performance of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ iterates using the refined SF-AdamW with $\beta_{1}\in\{0.1,0.5\}$ and $C=50$. Middle: Refined SF-AdamW with $\beta_{1}\in\{0.95,0.9\}$ and $C=200$ achieves improved performance over the best vanilla SF-AdamW run. Right: Under a large batch size (2M tokens), vanilla SF-AdamW with $\beta_{1}=0.98$ underperforms compared to AdamW with a cosine schedule, while the refined SF-AdamW—with only a sweep over $C=200$—matches its final performance.
@@ 6 Conclusion
While our findings highlight the potential of SF methods, several open questions remain. Our theoretical analysis relies on simplifying assumptions, and validating the central flow approximation in deep learning is a natural next step. Furthermore, extending the river-valley framework to analyze other modern optimizers—and exploring their integration with SF methods—offers a promising direction for further investigation. Lastly, our experiments are limited to small-scale language models due to computational constraints. Scaling these findings to larger models and longer training durations remains an important direction for future work.
Acknowledgements
This work was supported by a National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00421203).
@@ Appendix A Pseudocode for Schedule-Free AdamW
For completeness, we present the pseudocode for the Schedule-Free AdamW (SF-AdamW) algorithm in Algorithm 1, along with our proposed refinement in Algorithm 2. The refinement introduces a decoupling parameter $C$ to independently control the averaging window, addressing the coupling issue discussed in Section 5.
@@
[…]
@@ B.1 Language Model Experiments
Our main experiments use a 124M-parameter LLaMA-style decoder-only transformer with SwiGLU activations [Shazeer, 2020], RoPE embeddings [Su et al., 2024], RMSNorm [Zhang and Sennrich, 2019], and alternating attention/MLP blocks (12 layers, 12 attention heads, hidden dimension 768). Additional results in Appendix C verify our findings with a 124M-parameter GPT-2-style transformer. Both architectures are implemented in llm-baselines with standard design choices.
@@
Our main experiments use the 6B-token subset of the SlimPajama dataset [Soboleva et al., 2023], available on Hugging Face. [^1] https://huggingface.co/datasets/DKYoon/SlimPajama-6B We tokenize with the GPT-2 tokenizer [Radford et al., 2019], which has a vocabulary size of 50,304. In Appendix C, we also validate our main findings on the OpenWebText2 dataset [Gao et al., 2020], [^2] https://huggingface.co/datasets/segyges/OpenWebText2 using the same setup.
@@
We train models using AdamW and SF-AdamW, with a short warmup phase comprising 5% of total steps. For large-batch runs with cosine decay, the learning rate is annealed to 10% of its peak. Main experiments use a batch size of 1,024 sequences of context length 512 tokens (0.5M tokens total), trained for 5,000 steps, amounting to roughly 2.5B tokens (1$\sim$ 1$\times$ 1 × Chinchilla scale). Large-batch experiments use a 2M-token batch size and 2,500 steps ($\sim$5B tokens, or 2$\sim$ 2$\times$ 2 × Chinchilla scale) to evaluate efficiency in the overtrained regime. Validation is performed during training using 3,200 sequences of context length 512 tokens ($\sim$1.6M tokens) to compute validation loss (perplexity) curves. For computing EWA, we use a decay factor of 0.99 for all experiments.
@@
We fix the weight decay to 0.1 in all experiments. Gradient clipping is set to 1.0 for AdamW and disabled (0.0) for SF-AdamW. We perform sweeps over the learning rate, momentum parameters, and the decoupling parameter $C$ (for refined SF-AdamW). Full configurations are provided in Tables 1, 2, 3 and 4.
@@ B.2 Edge of Stability Experiments on CIFAR-10
We provide the experimental setup for Figure 6 in Section 4.3, where we study whether Schedule-Free methods operate at the Edge of Stability in a deep learning setting.
@@ Appendix C Additional Results
In this section, we present additional experiments on the OpenWebText2 dataset using a 124M-parameter GPT-2 style decoder-only transformer, replicating the setups from the main text. As summarized below, the results confirm that our main findings (Figures 2, 3 and 5) generalize across both datasets and architectures.
@@
  • Figure 2: As in the main experiments, we perform a grid search to determine the best hyperparameters for both AdamW and SF-AdamW, and evaluate whether LR decay or EWQ improves performance. In this setup, the best-performing configuration for AdamW is $(\beta_{1},\beta_{2})=(0.9,0.95)$ with LR 1e-3, while SF-AdamW achieves optimal results with $(\beta_{1},\beta_{2})=(0.95,0.99)$ and LR 2e-3. We observe that neither LR decay nor EWA improves performance for SF-AdamW, indicating that it already tracks the river closely (see Figure 9).
@@
  • Section 3: We also assess the behavior of SF-AdamW under suboptimal momentum configurations, using $\beta_{1}\in\{0.1,0.5\}$ while keeping all other settings fixed. For both values, we observe that applying a short LR decay phase using AdamW significantly improves performance, reducing the validation loss compared to the constant LR baseline. This result shows that SF-AdamW is sensitive to the choice of momentum, and suboptimal settings can hinder its ability to effectively follow the river (see Figure 10).
@@
  • Figure 5: Using the same runs from above, we compute the validation loss at the ${\mathbf{y}}_{t}$ iterates and their EWA. For suboptimal momentum settings ($\beta_{1}\in\{0.1,0.5\}$), the loss at ${\mathbf{y}}_{t}$ is consistently lower than that at ${\mathbf{x}}_{t}$, consistent with trends observed in both the toy model and SlimPajama experiments. Moreover, across all momentum values, applying EWA to ${\mathbf{y}}_{t}$ further improves performance, suggesting that ${\mathbf{y}}_{t}$ is more robust to suboptimal momentum and remains better aligned with the river trajectory (see Figure 11).
@@
  • Refined SF-AdamW: We evaluate the refined variant of SF-AdamW by sweeping over $C$ across multiple momentum settings. In the low-momentum regime ($\beta_{1}=0.5$), the refined method enables ${\mathbf{x}}_{t}$ to match or outperform ${\mathbf{y}}_{t}$, addressing the discrepancy observed in the vanilla formulation. In the best-performing vanilla configuration ($\beta_{1}=0.95$, LR 2e-3), setting $C=200$ yields further performance gains. Notably, comparable results can also be achieved under a suboptimal momentum setting ($\beta_{1}=0.9$) by choosing $C=50$, demonstrating improved robustness to hyperparameter choices. In the large-batch setting (2M-token batches), vanilla SF-AdamW with a constant LR ($(\beta_{1},\beta_{2})=(0.98,0.99)$, LR 2e-3) underperforms relative to AdamW with a cosine LR schedule ($(\beta_{1},\beta_{2})=(0.9,0.95)$, LR 5e-3). However, the refined variant with $C=500$ ($(\beta_{1},\beta_{2})=(0.98,0.99)$, LR 5e-3) successfully closes this performance gap. Together, this suggest that our refined method is robust to momentum and batch size scaling (see Figure 12).
@@
Figure 9: OpenWebText2 Experiment: SF-AdamW closely follows the river, unlike AdamW. Left, Middle: While AdamW benefits from linear LR decay and EWA, SF-AdamW shows no improvement from either. Right: A short decay phase of AdamW (with linear LR decay from 1e-4 to 0) leads to a sharp loss drop for AdamW, but has no effect when applied to the SF-AdamW trajectory—suggesting that SF-AdamW already tracks the river throughout training (Figure 2).
@@
Figure 10: OpenWebText2 Experiment: SF-AdamW with suboptimal momentum fails to follow the river. A short decay phase of AdamW applied to SF-AdamW checkpoints with $\beta_{1}\in\{0.1,0.5\}$ results in a sharp loss drop, unlike the case with $\beta_{1}=0.95$ (Section 3).
@@
Figure 11: OpenWebText2 Experiment: Performance of ${\mathbf{x}}_{t}$, ${\mathbf{y}}_{t}$, and the EWA of ${\mathbf{y}}_{t}$ under varying $\beta_{1}$. For suboptimal $\beta_{1}$, ${\mathbf{y}}_{t}$ outperforms ${\mathbf{x}}_{t}$, and across all momentum settings, the EWA of ${\mathbf{y}}_{t}$ achieves the lowest loss (Figure 5).

### v4

@@ 4.4 A Reformulation of Schedule-Free Optimizer
Motivated by Figures˜6 and 7, we now examine the dynamics of the ${\mathbf{y}}_{t}$ iterates. Following Defazio et al. (2024) and Morwani et al. (2025), define the momentum variable $\mathbf{m}_{t}\!\coloneqq\!\tfrac{\mathbf{x}_{t}-\mathbf{z}_{t+1}}{\gamma}$. The Schedule-Free update in (SF) is then equivalent to
@@
i.e. SF is a momentum-based optimizer update on $\mathbf{y}_{t}$. The full derivation is provided in Section˜D.2.
@@ 4.5 Central Flow Analysis
As $\beta_{1}$ increases, $\sigma^{2}$ decreases; thus, larger values of $\beta_{1}$ suppress oscillations along the hill directions, keeping the $\mathbf{y}_{t}$ iterates more closely aligned with the river—consistent with the empirical observation in Figure˜5. A complete derivation, including the central flow of SF-GD, is provided in Section˜D.3.
@@ 5 A Refined and Robust Schedule-Free Optimizer
Figure 8: Averaging weights in SF method. Smaller values of $\beta$ flatten the averaging weights $\{\alpha_{t}\}$.
While SF-AdamW achieves strong performance, it is highly sensitive to momentum hyperparameters and degrades under large batch sizes (Zhang et al., 2025; Morwani et al., 2025). Building on the insights from Section˜4, we revisit these issues and propose a refined variant that addresses them.
@@
When $\beta$ is small, the weights $\{\alpha_{t}\}$ become stretched, overemphasizing early iterates and preventing $\mathbf{x}_{t}$ from closely tracking the river, as shown in Figure˜8. This also explains Figure˜3, where we demonstrated that applying EWA to ${\mathbf{x}}_{t}$ offers no benefit: since ${\mathbf{x}}_{t}$ is already a weighted average of ${\mathbf{y}}_{t}$, further averaging merely flattens the weights and weakens alignment with the river.
More fundamentally, $\beta$ plays a dual role: it controls both the momentum update of ${\mathbf{y}}_{t}$ and the width of the averaging window for ${\mathbf{x}}_{t}$. The optimal value for each may differ, and this mismatch can hinder performance. In large-batch training, a narrower averaging window (i.e., larger $\beta$) is preferred to emphasize recent iterates. However, a large $\beta$ also slows the update of ${\mathbf{y}}_{t}$, as the $(1-\beta)\Delta_{t}$ term in (SFy) becomes small, reducing the influence of recent gradients.
@@
The full derivation is given in Section˜D.4, and pseudocode is provided in Algorithm˜2. As shown in Figures˜14 and 15, the weights $\{\alpha_{t}\}$ closely follow the theoretical approximation across different values of $\beta$, $C$, and $T$. Notably, both vanilla and refined SF use an averaging window that widens with $T$, unlike fixed-width schemes such as EWA.
This modification makes the averaging weights $\{\alpha_{t}\}$ depend solely on $C$, allowing $\beta$ to independently control the momentum on $\mathbf{y}_{t}$. In other words, $C$ decouples the momentum and averaging behavior. Below, we present preliminary experiments demonstrating the benefits of this refinement.
@@
We evaluate the performance of our refinement to SF-AdamW using the experimental setup as in Section˜3. Keeping all other hyperparameters fixed as the tuned vanilla configuration and varying only $C$ (with $C=1/(1-\beta_{1})$ recovering the original method), we observe:
  • Momentum robustness. For $\beta_{1}\in\{0.1,\,0.5\}$, the refined SF-AdamW allows ${\mathbf{x}}_{t}$ to match or outperform ${\mathbf{y}}_{t}$, in contrast to the underperformance of ${\mathbf{x}}_{t}$ reported in Figure˜6 (Figure˜9-left).
  • Improved best-case performance. In the best vanilla setup ($\beta_{1}=0.95$), setting $C=200$ leads to further reductions in validation loss; similar gains are observed with $\beta_{1}=0.9$ (Figure˜9-middle).
  • Large-batch setting. With 2M-token batches, vanilla SF-AdamW ($\beta_{1}=0.98$) lags behind AdamW with cosine schedule, whereas setting $C=200$ matches its final performance (Figure˜9-right).
To further examine the sensitivity to the new hyperparameter $C$, we conduct a set of controlled sweeps (Tables˜5 and 6 in Section˜C.1). Across both optimal and suboptimal $\beta_{1}$, we find that refined SF-AdamW consistently outperforms the vanilla baseline over a wide range of $C$, demonstraing its robustness to this parameter. These results show that decoupling momentum and averaging via introducing $C$ eliminates the tradeoff inherent in vanilla SF, yielding a more robust and scalable optimizer.
Figure 9: Refined SF-AdamW. Left: Performance of ${\mathbf{x}}_{t}$ and ${\mathbf{y}}_{t}$ iterates using the refined SF-AdamW with $\beta_{1}\in\{0.1,0.5\}$ and $C=50$. Middle: Refined SF-AdamW with $\beta_{1}\in\{0.95,0.9\}$ and $C=200$ achieves improved performance over the best vanilla SF-AdamW run. Right: Under a large batch size (2M tokens), vanilla SF-AdamW with $\beta_{1}=0.98$ underperforms compared to AdamW with a cosine schedule, while the refined SF-AdamW—with only a sweep over C=200C=200—matches its final performance.
@@ 6 Conclusion
While our findings highlight the potential of SF methods, several open questions remain. Our theoretical analysis relies on simplifying assumptions, and validating the central flow approximation in deep learning is a natural next step. Furthermore, extending the river-valley framework to analyze other modern optimizers, as well as exploring their integration with SF methods, is a promising direction for further investigation. Lastly, our experiments are limited to small-scale language models due to computational constraints. Scaling these findings to larger models and longer training durations remains an important direction for future work.
Acknowledgments and Disclosure of Funding
This work was supported by a National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No. RS-2024-00421203) and the InnoCORE program of the Ministry of Science and ICT (No. N10250156).
@@ Appendix A Pseudocode for Schedule-Free AdamW
For completeness, we present the pseudocode for the Schedule-Free AdamW (SF-AdamW) algorithm in Algorithm˜1, along with our proposed refinement in Algorithm˜2. The refinement introduces a decoupling parameter $C$ to independently control the averaging window, addressing the coupling issue discussed in Section˜5.
@@
11:  $c_{t+1}={\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\min}\left\{\frac{\gamma^{2}_{t}}{\sum^{t}_{i=1}\gamma^{2}_{i}}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}\cdot(1-\beta_{1})C,\;1}\right\}$
@@ B.1 Language Model Experiments
Our main experiments use a 124M-parameter LLaMA-style decoder-only transformer with SwiGLU activations [Shazeer, 2020], RoPE embeddings [Su et al., 2024], RMSNorm [Zhang and Sennrich, 2019], and alternating attention/MLP blocks (12 layers, 12 attention heads, hidden dimension 768). Additional results in Appendix˜C verify our findings with a 124M-parameter GPT-2-style transformer. Both architectures are implemented in llm-baselines with standard design choices.
@@
Our main experiments use the 6B-token subset of the SlimPajama dataset [Soboleva et al., 2023], available on Hugging Face. [^1] https://huggingface.co/datasets/DKYoon/SlimPajama-6B We tokenize with the GPT-2 tokenizer [Radford et al., 2019], which has a vocabulary size of 50,304. In Appendix˜C, we also validate our main findings on the OpenWebText2 dataset [Gao et al., 2020], [^2] https://huggingface.co/datasets/segyges/OpenWebText2 using the same setup.
@@
We train models using AdamW and SF-AdamW, with a short warmup phase comprising 5% of total steps. For large-batch runs with cosine decay, the learning rate is annealed to 10% of its peak. Main experiments use a batch size of 1,024 sequences of context length 512 tokens (0.5M tokens total), trained for 5,000 steps, amounting to roughly 2.5B tokens ($\sim 1\times$ Chinchilla scale). Large-batch experiments use a 2M-token batch size and 2,500 steps ($\sim$5B tokens, or $\sim 2\times$ Chinchilla scale) to evaluate efficiency in the overtrained regime. Validation is performed during training using 3,200 sequences of context length 512 tokens ($\sim$1.6M tokens) to compute validation loss (perplexity) curves. For computing EWA, we use a decay factor of 0.99 for all experiments.
@@
We fix the weight decay to 0.1 in all experiments. Gradient clipping is set to 1.0 for AdamW and disabled (0.0) for SF-AdamW. We perform sweeps over the learning rate, momentum parameters, and the decoupling parameter $C$ (for refined SF-AdamW). Full configurations are provided in Tables˜1, 2, 3 and 4.
@@ B.2 Edge of Stability Experiments on CIFAR-10
We provide the experimental setup for Figure˜7 in Section˜4.3, where we study whether Schedule-Free methods operate at the Edge of Stability in a deep learning setting.
@@ Appendix C Additional Results
In this section, we present additional experiments on the OpenWebText2 dataset using a 124M-parameter GPT-2 style decoder-only transformer, replicating the setups from the main text. As summarized below, the results confirm that our main findings (Figures˜3, 3 and 6) generalize across both datasets and architectures.
@@
  • Figure˜3: As in the main experiments, we perform a grid search to determine the best hyperparameters for both AdamW and SF-AdamW, and evaluate whether LR decay or EWQ improves performance. In this setup, the best-performing configuration for AdamW is $(\beta_{1},\beta_{2})=(0.9,0.95)$ with LR 1e-3, while SF-AdamW achieves optimal results with $(\beta_{1},\beta_{2})=(0.95,0.99)$ and LR 2e-3. We observe that neither LR decay nor EWA improves performance for SF-AdamW, indicating that it already tracks the river closely (see Figure˜10).
@@
  • Section˜3: We also assess the behavior of SF-AdamW under suboptimal momentum configurations, using $\beta_{1}\in\{0.1,0.5\}$ while keeping all other settings fixed. For both values, we observe that applying a short LR decay phase using AdamW significantly improves performance, reducing the validation loss compared to the constant LR baseline. This result shows that SF-AdamW is sensitive to the choice of momentum, and suboptimal settings can hinder its ability to effectively follow the river (see Figure˜11).
@@
  • Figure˜6: Using the same runs from above, we compute the validation loss at the ${\mathbf{y}}_{t}$ iterates and their EWA. For suboptimal momentum settings ($\beta_{1}\in\{0.1,0.5\}$), the loss at ${\mathbf{y}}_{t}$ is consistently lower than that at ${\mathbf{x}}_{t}$, consistent with trends observed in both the toy model and SlimPajama experiments. Moreover, across all momentum values, applying EWA to ${\mathbf{y}}_{t}$ further improves performance, suggesting that ${\mathbf{y}}_{t}$ is more robust to suboptimal momentum and remains better aligned with the river trajectory (see Figure˜12).
@@
  • Refined SF-AdamW: We evaluate the refined variant of SF-AdamW by sweeping over $C$ across multiple momentum settings. In the low-momentum regime ($\beta_{1}=0.5$), the refined method enables ${\mathbf{x}}_{t}$ to match or outperform ${\mathbf{y}}_{t}$, addressing the discrepancy observed in the vanilla formulation. In the best-performing vanilla configuration ($\beta_{1}=0.95$, LR 2e-3), setting $C=200$ yields further performance gains. Notably, comparable results can also be achieved under a suboptimal momentum setting ($\beta_{1}=0.9$) by choosing $C=50$, demonstrating improved robustness to hyperparameter choices. In the large-batch setting (2M-token batches), vanilla SF-AdamW with a constant LR ($(\beta_{1},\beta_{2})=(0.98,0.99)$, LR 2e-3) underperforms relative to AdamW with a cosine LR schedule ($(\beta_{1},\beta_{2})=(0.9,0.95)$, LR 5e-3). However, the refined variant with $C=500$ ($(\beta_{1},\beta_{2})=(0.98,0.99)$, LR 5e-3) successfully closes this performance gap. Together, this suggest that our refined method is robust to momentum and batch size scaling (see Figure˜13).
@@
Figure 10: OpenWebText2 Experiment: SF-AdamW closely follows the river, unlike AdamW. Left, Middle: While AdamW benefits from linear LR decay and EWA, SF-AdamW shows no improvement from either. Right: A short decay phase of AdamW (with linear LR decay from 1e-4 to 0) leads to a sharp loss drop for AdamW, but has no effect when applied to the SF-AdamW trajectory—suggesting that SF-AdamW already tracks the river throughout training (Figure˜3).
@@
Figure 11: OpenWebText2 Experiment: SF-AdamW with suboptimal momentum fails to follow the river. A short decay phase of AdamW applied to SF-AdamW checkpoints with $\beta_{1}\in\{0.1,0.5\}$ results in a sharp loss drop, unlike the case with $\beta_{1}=0.95$ (Section˜3).
@@
Figure 12: OpenWebText2 Experiment: Performance of ${\mathbf{x}}_{t}$, ${\mathbf{y}}_{t}$, and the EWA of ${\mathbf{y}}_{t}$ under varying $\beta_{1}$. For suboptimal $\beta_{1}$, ${\mathbf{y}}_{t}$ outperforms ${\mathbf{x}}_{t}$, and across all momentum settings, the EWA of ${\mathbf{y}}_{t}$ achieves the lowest loss (Figure˜6).
@@
C.1 Sensitivity to the Refinement Parameter $C$
To examine the sensitivity of refined SF-AdamW to the choice of the refinement parameter $C$, we conduct experiments on SlimPajama (0.5M and 2M token batch sizes). We sweep a range of $C$ values and report test perplexity over 9M held-out tokens. The results are summarized in Table 5 and Table 6.
Our results demonstrate that refined SF-AdamW is consistently robust to the choice of $C$. Across a broad range of settings, it outperforms vanilla SF-AdamW (i.e., $C=1/(1-\beta_{1})$). For instance, in Table 5, values such as $C\in\{50,100,200,500\}$ improve performance for $\beta_{1}=0.9,0.95$, and $C=50,100$ improve performance across all momentum values. Similar trends hold in Table 6, with performance gains persisting even at large $C$. These results indicate that refined SF-AdamW remains effective without requiring careful tuning of $C$.
$\beta_{1}$ Vanilla $C=5$ 10 20 50 100 200 500
0.1 67.20 37.06 35.80 36.62 36.15 37.27
0.5 41.01 29.32 30.87 29.96 29.57
0.9 27.70 27.70 23.97 24.64 24.93 25.11
0.95 25.12 25.12 23.98 23.60 24.37 24.83
Table 5: Refined SF-AdamW on SlimPajama with 0.5M-token batch size. Test perplexity (on 9M held-out tokens) under varying $C$ values. LR = 2e-3, $\beta_{2}=0.99$.
$\beta_{1}$ Vanilla $C=5$ 10 20 50 100 200 500 1000
0.1 110.8 54.80 47.37 44.31 43.80
0.5 54.24 47.80 38.42 38.86 38.49 42.97
0.9 31.34 31.34 29.86 27.68 28.16 27.02 27.88
0.95 27.29 27.29 25.77 25.45 25.77 27.31
0.98 26.09 30.23 26.09 25.49 24.51 23.95 23.88
Table 6: Refined SF-AdamW on SlimPajama with 2M-token batch size. Test perplexity (on 9M held-out tokens) under varying $C$ values. LR = 2e-3, $\beta_{2}=0.99$.